# Tiny models from tiny data: Textual and null-text inversion for few-shot distillation

### v1

@@ 5.4 Computational performance
Measured using one NVidia A40 GPU, the total time required for specialization to 5 novel classes from one miniImageNet episode is 6.41 hours, with 5.59h for generating 20k images (4k per class) using TINT and 0.82h for distillation. In contrast, NAO+SeedSelect [32] requires 47.5h GPU hours for generating 5k images (1k per class). This equals 1.0s per generated image for our method compared to 34.2s for NAO+SeedSelect. See Appendix A.4 for a detailed breakdown.
@@ 6 Discussion
Fairness of using extra data Few-shot classification benchmarks have traditionally been used in a strict setting, without allowing extra data. Nowadays, this strict view if often relaxed, and there are several published few-shot results relying on extra data [8, 42, 32]. Other benchmarks, e.g. regular ImageNet classification, have undergone similar transitions, where most top results now rely on extra training data. As discussed in Section 5.3, we find it reasonable to compare results with/without extra data as long as the use of extra data is clearly disclosed. We fully acknowledge that the strict setting is more challenging, but note that many real-world applications have no rule against extra data. We believe that practitioners will find it useful to see how large accuracy gains that are achievable by allowing extra data in the training.
Outlook Image-generating diffusion models improve in a rapid pace. In our pipeline, it is straight-forward to replace the diffusion model with an improved one. As such models become even more general, we expect more and more applications to be fully solvable using our method, even when small, highly efficient final models are required and only a tiny set of application-specific data is available. We hope that this approach will be adapted by practitioners, and that our work will inspire the use of generative models in a wider range of data-constrained problems.
@@ Acknowledgements
This work was partially funded by Pixelwise AB and was supported by the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation. The experiments were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ A.1 Pseudo-code
     $\mathbf{z}_{n,0,1:K}\leftarrow\textrm{vqvae\_encode}(\mathbf{x}^{\textrm{(s)}}_{n,1:K})$
     $\mathbf{v}_{n}\leftarrow\textrm{textinv}(\mathbf{z}_{n,0,1:K})$
@@
         $(\mathbf{u}_{n,1:t,k},\mathbf{z}_{n,T,k})\leftarrow\textrm{nulltextinv}(\mathbf{z}_{n,0,k})$
@@
     return $(\mathbf{v}_{n},\mathbf{u}_{n,1:t,1:K},\mathbf{z}_{n,T,1:K})$
@@
procedure GenGenerate($\mathbf{v}_{n},\mathbf{u}_{n,1:t,1:K},\mathbf{z}_{n,T,1:K}$)
@@
     $\mathbf{z}\leftarrow(1-\alpha)\mathbf{z}_{n,T,k}+\alpha\mathbf{n}$
     $\mathbf{\tilde{z}}\leftarrow\mathbf{z}\|\mathbf{z}\|^{-1}\big{(}(1-\alpha)\|\mathbf{z}_{n,T,k}\|+\alpha\|\mathbf{n}\|\big{)}$
     $\mathbf{z}_{0}\leftarrow G(\mathbf{\tilde{z}}\>;\mathbf{v}_{n},\mathbf{u}_{n,1:t,k})$
@@ A.2 Implementation details
The code required to run all experiments was implemented in PyTorch [80] using the Diffusers library [81] with Stable Diffusion 1.5. The overall experiment framework was implemented by us from scratch, but includes code parts from prior work [24, 23, 8, 31, 32]. Full source code will be made available at https://github.com/pixwse/tiny2.
A.3 Estimator variance
This section presents more details on the analysis of the accuracy estimation, including a proof of Theorem 1.
Proof of Theorem 1 Computing the estimated accuracy $\tilde{a}_{p}$ over an episode $p$ involves classifying $Q$ independent query examples using a fixed classifier. This can be modeled as performing $Q$ Bernoulli trials with success rate $a_{p}$. $\tilde{a}_{p}$ is then the mean of $Q$ Bernoulli-distributed variables, which follows a binomial distribution with
@@ A.3 Estimator variance
$\mathbb{V}[\tilde{a}_{p}]$ $=\frac{1}{Q}a_{p}(1-a_{p}).$ (8)
Considering that episodes are drawn randomly and that each episode may have a different true accuracy $a_{p}$, we model each $a_{p}$ as a random variable with $\mathbb{E}[a_{p}]=a$ and $\mathbb{V}[a_{p}]=\sigma_{a}^{2}$. Since episodes are drawn randomly and independently, all $a_{p}$ are i.i.d.
For the expectation, we simply have $\mathbb{E}[\tilde{a}_{p}|a_{p}]=\mathbb{E}[a_{p}]=a$. To determine $\mathbb{V}[\tilde{a}_{p}]$, we need to consider two sources of variance: due to estimating each $\tilde{a}_{p}$ from a finite query set (intra-episode variance) and due to each episode having a different true accuracy (inter-episode variance) based on the intrinsic difficulty of each episode. Starting at the law of total variance (Eve’s law) [82], we can write
$\mathbb{V}[\tilde{a}_{p}]$ $=\mathbb{E}\big{[}\mathbb{V}[\tilde{a}_{p}|a_{p}]\big{]}+\mathbb{V}\big{[}\mathbb{E}[\tilde{a}_{p}|a_{p}]\big{]}=$ (9)
$=\mathbb{E}\left[\frac{1}{Q}a_{p}(1-a_{p})\right]+\mathbb{V}[a_{p}]=$ (10)
@@
From the well-known variance formula $\mathbb{V}[x]=\mathbb{E}[x^{2}]-\mathbb{E}[x]^{2}$, we get $\mathbb{E}[a_{p}^{2}]=\mathbb{V}[a_{p}]+\mathbb{E}[a_{p}]^{2}=\sigma_{a}^{2}+a^{2}$ and arrive at
$\mathbb{V}[\tilde{a}_{p}]$ $=\frac{1}{Q}(a-\sigma_{a}^{2}-a^{2})+\sigma_{a}^{2}=$ (12)
@@
$\mathbb{V}[\tilde{a}]=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\left(1-\frac{1}{Q}\right)\sigma_{a}^{2}\right)$ (14)
@@
Discussion and usage example Theorem 1 allows us to study the expected impact of varying $Q$ and $P$. Note that for moderately large $Q$, Eq. 14 can be approximated as
$\mathbb{V}[\tilde{a}]=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\sigma_{a}^{2}\right).$ (15)
As $Q$ increases, $\mathbb{V}[\tilde{a}]$ approaches $P^{-1}\sigma_{a}^{2}$, representing the case where the accuracy of each episode is estimated perfectly, and only the inter-episode variance remains.
Eq. 14 (or Eq. 15) can be used for experiment planning, assessing the expected variance of an evaluation using a given $P$ and $Q$, or giving guidance around the number of evaluation episodes required for demonstrating a certain improvement with a desired statistical significance. As a concrete example, for the LabelHalluc method [31], $a\approx 0.87,\sigma\approx 0.05$, giving an estimator variance of 7e-6 over 600 episodes using 25 query examples (5 per class). In preliminary experiments, the accuracy of our method was $a=0.93$ with $\sigma=0.028$. Increasing $P$ to include all remaining examples ($595N$), asserting that we want a similar estimator variance as the original LabelHalluc method and solving for $P$ gives us $P=118$. We therefore settled for 120 evaluation episodes as a reasonable trade-off between computational requirements and confidence. As can be seen in Table 3, our 95% confidence interval for ResNet12 on miniImageNet is indeed similar to that of the LabelHalluc method (0.50 vs 0.48), showing that the practical results are consistent with the theory.
@@ B.1 Models and pre-training
Students The network used in the original ProtoNet method [64] was dubbed Conv4 in later works [37, 83, 84]. It is made up of 4 blocks, each consisting of a 3$\times$3 convolution, batchnorm, ReLU and 2$\times$2 max pooling. The first block increases the channel count to 64, which is kept throughout the 4 blocks. The ResNet12 is another popular choice [31, 85, 65, 37]. With 8M parameters, it is significantly larger than Conv4, but still one of the smaller networks used in prior art. It consists of 4 residual modules, each one consisting of 3 blocks. One block consists of a 3$\times$3 convolution, bachnorm and ReLU. A skip connection connects the input of the first block to the input of the ReLU of the last block. The scale is decreased by a 2$\times$2 max pooling at the end of each module. The first convolution of each module increases the channel count to 64, 128, 256 and 512 channels respectively. Both our implementation and pre-training using IER [86] directly follows prior work [31].
Teacher The teacher backbone used DINO ViT-B/16 with pre-trained weights. The teacher pretraining largely followed the original P$>$M$>$F paper [8], with some minor variations due to implementation details. We used the ADAM optimizer with initial learning rate 3e-3, lowered by a cosine schedule over 100 epochs. Each epoch consisted of 400 episodes of 5 classes with 5 support examples and 15 query examples per class. The optimizer used a momentum term of 0.9 and weight decay 1e-6. No test-time fine tuning was performed for the teacher, since this is only important when evaluating cross-domain performance.
@@ B.2 Synthetic data generation
Diffusion model For data generation, we used Stable Diffusion 1.5, with the DDIM scheduler running over 50 time steps. The textual inversion was implemented largely following the original paper [23], with 5k optimization iterations, using the ADAM optimizer with fixed learning rate 1e-3. The null-text inversion was run using 10 iterations per timestep, using the ADAM optimizer with a learning rate varying over timesteps according to the original implementation [24].
Resolution mismatch  To handle the resolution mismatch between mini-imagenet and the pre-trained Stable Diffusion model, the 84$\times$84 images were first upsampled to 96$\times$96, then upscaled with super resolution using HAT [33] to 384$\times$384, and then upsampled again to 512$\times$512. Bilinear interpolation was used in the two upsampling steps. Some hand-picked examples of generated images with and without super-resolution are shown in Figure 6. Note that some images in the top row are completely degenerate (1 and 5), and that there are traces of a grid-like structure also in the other images (carpets in image 2-4), showing why the super resolution step is needed.
@@ B.3 Datasets
For miniImageNet, we used the common train-val-test split from [21]. For CUB, we used the common 100/50/50 split (randomly selected) and with images cropped and downsampled to 84x84 resolution. For CIFAR-FS, we used the original splits from [70]. The exact episodes used are provided with the source code. We encourage follow-up work to use the same episodes when comparing with our work, but also to run a final validation on a new set of random episodes to avoid over-engineering hyper-parameters for one specific fixed episode set.
@@ B.4 Distillation
Similar to prior work  [31], we used an SGD optimizer with momentum 0.9 and weight decay 5e-4. We adjusted the batch size to 150 to enable constructing a batch from equal parts of 1-3 data sources. The learning rate was set to 0.02 (adjusted slightly due to the change of batch size). For the longer runs, it was reduced by a factor 0.1 after 5k iterations. The loss was computed using cross-entropy for the support examples and KL-divergence for the base and synthetic data sources.
@@ B.5 LAION data mining details
The LAION dataset consists of 5.85 billion image-text pairs. Working with the entire dataset is unpractical, as it weighs in at 240Tb (and already the metadata requires 800Gb). Furthermore, many images in LAION are of rather poor quality. In Stable Diffusion, the final iterations of the training is performed a subsets of LAION with extra high aesthetic quality, assessed by an automated method. As a practical compromise, we use the subset improved aesthetics 6+, consisting of 12M images.
@@ B.6 Statistical significance
Theorem 1 was used to guide experiment planning, arriving at 120 episodes but evaluating using all remaining query examples per episode as a reasonable trade-off between significance and computation time (see Section A.3). The final 95% confidence intervals reported in Table 3 are computed from the individual episode accuracies using the Students-t distribution as usual, directly following prior work [31].
B.7 Conv4 ablations
In addition to the experiments shown in Section 5.2, we run a few selected ablations/comparisons using the Conv4 backbone on the same 5 fixed episodes of the validation split of miniImageNet. The results are shown in in Table 6. Note that no hyperparameter tuning was applied here, so the results can likely be improved. However, the results show that it is at least not straight-forward to reach good performance using the competing NAO+SeedSelect method [32]. The results also show that TINT significantly beats plain textual inversion also for this tiny student model.
Table 5: Results for the Conv4 backbone on the 5 fixed episodes, miniImageNet validation split.
Method Syn teacher Acc
Plain TI (4k) PMF 70.4
TINT (4k) PMF 81.7
NAO+SS [32] (1k) None 67.3
NAO+SS [32] (1k) PMF 64.0
TINT (1k) PMF 80.9
Table 6: FID measure (64 features) for different synthetic data generation choices.
Method FID $\downarrow$
NAO+SeedSelect [32] 13.0
Plain TI 8.0
TI + superres + limited t range 4.0
Full TINT 2.7
B.8 FID experiments
One way of evaluating generative models is using the FID measure [68]. The original measure used the 2048-dimensional feature vector from an Inception V3 network. In our case, we want to compare generated images with real images drawn from miniImageNet, where there are only 600 examples per class. Estimating a $2048\times 2048$ covariance matrix from only a few hundred images is numerically problematic. Instead, following prior work [32], we use a feature space of only 64 dimensions, using pooled features from an earlier layer of the Inception V3 model. We generated 300 random images per class, compared to 300 randomly drawn real images of the same class, repeated the process for 5 episodes (with 5 classes per episode) and computed the average FID score. The results are shown in Table 6. Eventhough this is a rather coarse measure, it shows that the TINT in some respect generates images that follow the novel class image distribution better than competing methods.
Appendix C External assets
This section gives an overview of external assets, download links and licenses. All assets that have no explicit license were made publicly available together with wording encouraging research use.
Appendix D Generated image examples

### v2

@@ 5.5 Computational performance
Measured using one NVidia A40 GPU, the total time required for specialization to 5 novel classes from one miniImageNet episode is 6.41h, with 5.59h for generating 20k images (4k per class) using TINT and 0.82h for distillation. In contrast, NAO+SeedSelect (Samuel et al., 2024a) requires 47.5h GPU hours for generating 5k images (1k per class). This equals 1.0s per generated image for our method compared to 34.2s for NAO+SeedSelect. See Appendix A.4 for a detailed breakdown.
@@ 6 Discussion
Fairness of using extra data Few-shot classification benchmarks have traditionally been used in a strict setting, without allowing extra data. Nowadays, this strict view if often relaxed, and there are several published few-shot results relying on extra data Hu et al. (2022); Fifty et al. (2024); Samuel et al. (2024a). Other benchmarks, e.g. regular ImageNet classification, have undergone similar transitions, where most top results now rely on extra training data. As discussed in Section 5.4, we find it reasonable to compare results with/without extra data as long as the use of extra data is clearly disclosed. We fully acknowledge that the strict setting is more challenging, but note that many real-world applications have no rule against extra data. We believe that practitioners will find it useful to see how large accuracy gains that are achievable by allowing extra data in the training.
Outlook Image-generating diffusion models improve in a rapid pace. In our pipeline, it is straight-forward to replace the diffusion model with an improved one. As such models become even more general, we expect more and more applications to be fully solvable using our method, even when small, highly efficient final models are required and only a tiny set of application-specific data is available.
Generative distillation is not restricted to classification. Similar methods could be used for e.g. segmentation by replacing the teacher with a strong (but slow) few-shot segmentation engine. Furthermore, if other information than text is suitable for defining classes in a specific application, the conditional text input could be replaced with any other conditioning mechanism. For example, for medical applications, we could use a generative model conditioned on e.g. tabular medical data. The textual inversion would then rather be a medical data inversion, but the method formulation will still apply.
We hope that this work has shown the potential of using generative models in few-shot distillation, and that it will inspire more research in this direction for a wider range of label-constrained problems.
@@ Acknowledgements
The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
@@ A.1 Pseudo-code
     $\mathbf{z}_{0,n,1:K}\leftarrow\textrm{vqvae\_encode}(\mathbf{x}^{\textrm{(s)}}_{n,1:K})$
     $\mathbf{v}_{n}\leftarrow\textrm{textinv}(\mathbf{z}_{0,n,1:K})$
@@
         $(\mathbf{u}_{1:t,n,k},\mathbf{z}_{T,n,k})\leftarrow\textrm{nulltextinv}(\mathbf{z}_{0,n,k})$
@@
     return $(\mathbf{v}_{n},\mathbf{u}_{1:t,n,1:K},\mathbf{z}_{T,n,1:K})$
@@
procedure GenGenerate($\mathbf{v}_{n},\mathbf{u}_{1:t,n,1:K},\mathbf{z}_{T,n,1:K}$)
@@
     $\mathbf{z}^{\prime}\leftarrow(1-\alpha)\mathbf{z}_{T,n,k}+\alpha\mathbf{n}$
     $\mathbf{\tilde{z}}\leftarrow\mathbf{z}^{\prime}\|\mathbf{z}^{\prime}\|^{-1}\big{(}(1-\alpha)\|\mathbf{z}_{n,T,k}\|+\alpha\|\mathbf{n}\|\big{)}$
     $\mathbf{z}_{0}\leftarrow G(\mathbf{\tilde{z}}\>;\mathbf{v}_{n},\mathbf{u}_{1:t,n,k})$
@@ A.2 Implementation details
The code required to run all experiments was implemented in PyTorch (Paszke et al., 2019) using the Diffusers library (Face, 2014) with Stable Diffusion 1.5. The overall experiment framework was implemented by us from scratch, but includes code parts from prior work (Mokady et al., 2023; Gal et al., 2022; Hu et al., 2022; Jian & Torresani, 2022; Samuel et al., 2024a). Source code is available at https://github.com/pixwse/tiny2.
A.3 Proof of Theorem 1
Computing the estimated accuracy $\tilde{a}_{p}$ over an episode $p$ involves classifying $Q$ independent query examples using a fixed classifier. This can be modeled as performing $Q$ Bernoulli trials with success rate $a_{p}$. $\tilde{a}_{p}$ is then the mean of $Q$ Bernoulli-distributed variables, which follows a binomial distribution with
@@ A.3 Proof of Theorem 1
$\mathrm{Var}[\tilde{a}_{p}]$ $=\frac{1}{Q}a_{p}(1-a_{p}).$ (9)
Considering that episodes are drawn randomly and that each episode may have a different true accuracy $a_{p}$, we model each $a_{p}$ as a random variable with $\mathbb{E}[a_{p}]=a$ and $\mathrm{Var}[a_{p}]=\sigma_{a}^{2}$. Since episodes are drawn randomly and independently, all $a_{p}$ are i.i.d.
For the expectation, we simply have $\mathbb{E}[\tilde{a}_{p}|a_{p}]=\mathbb{E}[a_{p}]=a$. To determine $\mathrm{Var}[\tilde{a}_{p}]$, we need to consider two sources of variance: due to estimating each $\tilde{a}_{p}$ from a finite query set (intra-episode variance) and due to each episode having a different true accuracy (inter-episode variance) based on the intrinsic difficulty of each episode. Starting at the law of total variance (Eve’s law) (Blitzstein & Hwang, 2019), we can write
$\mathrm{Var}[\tilde{a}_{p}]$ $=\mathbb{E}\big{[}\mathrm{Var}[\tilde{a}_{p}|a_{p}]\big{]}+\mathrm{Var}\big{[}\mathbb{E}[\tilde{a}_{p}|a_{p}]\big{]}=$ (10)
$=\mathbb{E}\left[\frac{1}{Q}a_{p}(1-a_{p})\right]+\mathrm{Var}[a_{p}]=$ (11)
@@
From the well-known variance formula $\mathrm{Var}[x]=\mathbb{E}[x^{2}]-\mathbb{E}[x]^{2}$, we get $\mathbb{E}[a_{p}^{2}]=\mathrm{Var}[a_{p}]+\mathbb{E}[a_{p}]^{2}=\sigma_{a}^{2}+a^{2}$ and arrive at
$\mathrm{Var}[\tilde{a}_{p}]$ $=\frac{1}{Q}(a-\sigma_{a}^{2}-a^{2})+\sigma_{a}^{2}=$ (13)
@@
$\mathrm{Var}[\tilde{a}]=\frac{1}{P}\left(\frac{1}{Q}a(1-a)+\left(1-\frac{1}{Q}\right)\sigma_{a}^{2}\right)$ (15)
@@ B.1 Models and pre-training
Students The network used in the original ProtoNet method (Snell et al., 2017) was dubbed Conv4 in later works (Ye et al., 2020; Afrasiyabi et al., 2020; Gidaris et al., 2019). It is made up of 4 blocks, each consisting of a 3$\times$3 convolution, batchnorm, ReLU and 2$\times$2 max pooling. The first block increases the channel count to 64, which is kept throughout the 4 blocks. The ResNet12 is another popular choice (Jian & Torresani, 2022; Lee et al., 2019; Xie et al., 2022; Ye et al., 2020). With 8M parameters, it is significantly larger than Conv4, but still one of the smaller networks used in prior art. It consists of 4 residual modules, each one consisting of 3 blocks. One block consists of a 3$\times$3 convolution, bachnorm and ReLU. A skip connection connects the input of the first block to the input of the ReLU of the last block. The scale is decreased by a 2$\times$2 max pooling at the end of each module. The first convolution of each module increases the channel count to 64, 128, 256 and 512 channels respectively. Both our implementation and pre-training using IER (Rizve et al., 2021) directly follows Jian & Torresani (2022).
Teacher The teacher backbone used DINO ViT-B/16 with pre-trained weights. The teacher pretraining largely followed the original P$>$M$>$F paper (Hu et al., 2022), with some minor variations due to implementation details. We used the ADAM optimizer with initial learning rate 3e-3, lowered by a cosine schedule over 100 epochs. Each epoch consisted of 400 episodes of 5 classes with 5 support examples and 15 query examples per class. The optimizer used a momentum term of 0.9 and weight decay 1e-6. No test-time fine tuning was performed for the teacher, since this is only important when evaluating cross-domain performance.
@@ B.2 Synthetic data generation
Diffusion model For data generation, we used Stable Diffusion 1.5, with the DDIM scheduler running over 50 time steps. The textual inversion was implemented largely following the original paper (Gal et al., 2022), with 5k optimization iterations, using the ADAM optimizer with fixed learning rate 1e-3. The null-text inversion was run using 10 iterations per timestep, using the ADAM optimizer with a learning rate varying over timesteps according to the original implementation (Mokady et al., 2023).
Resolution mismatch  To handle the resolution mismatch between mini-imagenet and the pre-trained Stable Diffusion model, the 84$\times$84 images were first upsampled to 96$\times$96, then upscaled with super resolution using HAT (Chen et al., 2023) to 384$\times$384, and then upsampled again to 512$\times$512. Bilinear interpolation was used in the two upsampling steps. Some hand-picked examples of generated images with and without super-resolution are shown in Figure 7. Note that some images in the top row are completely degenerate (1 and 5), and that there are traces of a grid-like structure also in the other images (carpets in image 2-4), showing why the super resolution step is needed.
@@ B.3 Datasets
For miniImageNet, we used the common train-val-test split from Ravi & Larochelle (2016). For CUB, we used the common 100/50/50 split (randomly selected) and with images cropped and downsampled to 84x84 resolution. For CIFAR-FS, we used the original splits from Bertinetto et al. (2018). The exact episodes used are provided with the source code. We encourage follow-up work to use the same episodes when comparing with our work, but also to run a final validation on a new set of random episodes to avoid over-engineering hyper-parameters for one specific fixed episode set.
@@ B.4 Distillation
Similar to prior work (Jian & Torresani, 2022), we used an SGD optimizer with momentum 0.9 and weight decay 5e-4. We adjusted the batch size to 150 to enable constructing a batch from equal parts of 1-3 data sources. The learning rate was set to 0.02 (adjusted slightly due to the change of batch size). For the longer runs, it was reduced by a factor 0.1 after 5k iterations. The loss was computed using cross-entropy for the support examples and KL-divergence for the base and synthetic data sources.
@@ B.5 LAION data mining details
The LAION dataset consists of 5.85 billion image-text pairs. Working with the entire dataset is unpractical, as it weighs in at 240Tb (and already the metadata requires 800Gb). Furthermore, many images in LAION are of rather poor quality. In Stable Diffusion, the final iterations of the training is performed on subsets of LAION with extra high aesthetic quality, assessed by an automated method. As a practical compromise, we use the subset improved aesthetics 6+, consisting of 12M images.
@@ B.6 Statistical significance
Theorem 1 was used to guide experiment planning, arriving at 120 episodes but evaluating using all remaining query examples per episode as a reasonable trade-off between significance and computation time (see Section A.3). The final 95% confidence intervals reported in Table 5 are computed from the individual episode accuracies using the Students-t distribution as usual, directly following Jian & Torresani (2022).
Appendix C Generated image examples
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption